OpenStudy (anonymous):

Please help!!! :) partial fraction of 3/ x^2-3x

OpenStudy (anonymous):

Start by factoring the denominator. \[3/((x)(x-3))\] This one is relatively simple to break up; to linear factors. \[A/x+B(x-3)=3/(x^2-3x)\] Now we clear the denominators.... \[A(x-3)+B(x)=3\] Using this, we generate a system of equations by equating the coefficients A+B=0x---(this equation corresponds to the X coefficients) -3A+0B=3----(this equation corresponds to the constants) If you're completely unsure of how to start the problem, your best bet is checking a textbook; I kind of blew through an explanation; hopefully it jogged your memory if you've seen it before.

OpenStudy (anonymous):

Also, from the system of equations, we know that A=-1 and B=1 This means that our final answer(our partial fraction decomposition) is \[-1/x+1/(x-3)\]

OpenStudy (anonymous):

thank you so so so much!!! :)